OpenStudy (anonymous):

PLEASEE HELPP!!! An angle measures 38° more than the measure of a supplementary angle. What is the measure of each angle? you need to answers

OpenStudy (anonymous):

supplementary means 180 degrees (2 angles sum to total 180)

OpenStudy (anonymous):

Two angles are supplementary if they add up to 180. If you have 2 angles adding to 180 and the angles are called x and y, we can say: x + y = 180 We can say that: y = 180 - x We also know that: y = x - 38 So, we can set: x - 38 = 180 - x And now you just have to solve for "x"

OpenStudy (anonymous):

ok hold on 1 second

OpenStudy (anonymous):

Once you get x, you have 180 - x for the other angle.

OpenStudy (anonymous):

ok i think i did something wrong i got a decimal

OpenStudy (anonymous):

ya i have no clue how to get x it keeps getting up to a decimal

OpenStudy (anonymous):

ok i got 109

OpenStudy (anonymous):

I think you need to look at this like this... x + y = 180 x = y + 38

OpenStudy (anonymous):

You have to get the "x" terms on one side and then get the numerics on the other. First step: add "x" to both sides. Second step: add 38 to both sides. Third step: divide both sides by the coefficient on "x". Fourth step: the other angle is 180 - x.

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

And now, you get 180 - 109 for the other.

OpenStudy (anonymous):

substitue and get y + 38 + y = 180 2 y + 38 = 180 2y = 180-38 2y = 142 y = 71 thus leaving x + y = 180 x + 71 = 180 x = 180-71 x = 109

OpenStudy (anonymous):

Now you have both angles and you're done.

OpenStudy (anonymous):

71 ya!!!! THANKS SO MUCH YOU HAVE BEEN VERY HELPFUL!!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @TeamJakeRosati_06

OpenStudy (anonymous):

your welcome and thanks :)